# Token-Domain Multiple Access: Exploiting Semantic Orthogonality for Collision Mitigation

### v1

@@ II-B Evolution of Multiple Access in 6G
Advancements in multiple access towards 6G focus on efficient random access for numerous uncoordinated devices with sporadic activity and short data packets. Key 6G technologies like NOMA and grant-free random access (GFRA) are expected to meet these demands [28, 29]. For example, in GFRA, after receiving a beacon signal from the base station (BS), devices transmit their unique preambles over the same time-frequency resources. The BS then performs device activity detection, channel state information (CSI) estimation, and data detection. This protocol is commonly utilized in studies such as [30, 31, 32]. To accommodate a larger number of devices, unsourced massive access (UMA) introduces a shared preamble codebook, where messages are represented as indices of codewords within the codebook. The BS focuses on decoding the transmitted codewords from the shared codebook, without necessarily associating them to their transmitters [33, 34, 35].
@@ IV-A2 Token Modulation
Assuming that the same tokenizer is shared among the devices, we propose to use a shared modulation codebook for multiple access, denoted as ${\bf U}\in\mathbb{C}^{L\times Q}$, where $L$ denotes the length of each codeword. For simplicity, we consider the columns of ${\bf U}$ are orthonormal vectors, i.e., ${\bf U}^{H}{\bf U}=\mathbf{I}_{Q}$. We use $g(\cdot)$ to represent the token modulation process; hence, the output is denoted by ${\bf X}_{k}={\bf U}{\bf B}_{k}=g({\bf B}_{k})\in\mathbb{C}^{L\times N}$, where each token is modulated to a modulation codeword. We denote the $n$-th column of ${\bf X}_{k}$ as ${\bf x}_{k,n}={\bf U}{\bf b}_{k,n}\in\mathbb{C}^{L}$. The advantages of the proposed token modulation are mainly two fold:
  • Shared modulation codebook enables uncoordinated random access of massive number of devices, which reduces the access latency and is easy for implementation.
  • The basic communication unit changes from the conventional “bits” to “tokens”, which facilitates applying transformer-based token processing and MLLM s at the receiver.
@@ IV-B Multiple Access Signal Model
$={\bf U}\sum\limits_{k\in[K]}{\bf b}_{k,n}\left({\bf h}_{k}\right)^{T}+{\bf Z}_{n}={\bf U}{\bf H}_{n}+{\bf Z}_{n},$ (1)
where ${\bf h}_{k}\in\mathbb{C}^{M}$ denotes the channel vector between the $k$-th device and the BS, ${\bf H}_{n}=\sum\limits_{k\in[K]}{\bf b}_{k,n}\left({\bf h}_{k}\right)^{T}\in\mathbb{C}^{Q\times M}$ is termed as an equivalent channel matrix, and the noise matrix ${\bf Z}_{n}$ follows an independent and identically distributed (i.i.d.) complex Gaussian distribution with zero mean and variance $\sigma^{2}{\bf I}_{n}$. Without loss of generality, the channel vector ${\bf h}_{k}$, $\forall k$, is assumed to remain the same throughout the $N$ time slots.
@@ V-C Token Assignment
Here, the CSI acts as a matching anchor that links detected tokens to their respective devices. We assume the CSI of $K$-devices are known, denoted by $\{{\bf h}_{1},{\bf h}_{2},...,{\bf h}_{K}\}$. We initialize the estimated token sequences as ${\bf c}_{k}={\bf 0}_{N}$, $\forall k\in[K]$. Note that the elements of ${\bf c}_{k}$ shall be integers within $[Q]$ to represent tokens, which is equivalent to the one-hot vector representation introduced in Section IV-A. The token assignment process is described as follows.
  • Initial Assignment: We define a residual token set $\mathcal{P}_{n}^{r}$, which is initialized as $\mathcal{P}_{n}^{r}=\mathcal{P}_{n}$, $\forall n\in[N]$. We iterate $k$ from $1$ to $K$, in each iteration, we first find the token that minimize the Euclidean distance:
@@
    where $\|\cdot\|_{2}$ denotes the $l_{2}$ norm. Next, if and only if $\|{\bf h}_{k}^{0}-{\bf h}_{\phi^{\star}_{n},n}\|_{2}^{2}/M<T_{h}$, we assign token $\phi^{\star}_{n}$ to $[{\bf c}_{k}]_{n}$, and then update $\mathcal{P}_{n}^{r}$ by deleting $\phi^{*}_{n}$; Otherwise, we assign “[MASK]” to $[{\bf c}_{k}]_{n}$.
  • Fine-Grained Update: After the initial assignment, if $\mathcal{P}_{n}^{r}=\emptyset$, there is no token collision in this time slot; if $|\mathcal{P}_{n}^{r}|=1$, we assign the only residual token in $\mathcal{P}_{n}^{r}$ to the “[MASK]” positions of $[{\bf c}_{k}]_{n}$, $\forall k\in[K]$; If $|\mathcal{P}_{n}^{r}|>1$, the elements of this residual token set serve as candidate tokens to assist the following steps.
@@ VI-A3 Token Communication Metrics
We define token error rate (TER) as $\text{TER}=\dfrac{1}{2NK}\sum_{k=1}^{K}\|\widehat{\bf B}_{k}-{\bf B}_{k}\|_{0}.$ It is worth emphasizing that TER is a general metric for ToDMA without specific data modalities. To evaluate the image transmission performance, we consider the well-known PSNR and LPIPS to measure the distortion and perception qualities, respectively.
@@ VI-B Latency Comparisons
To calculate the communication latency, we consider an OFDM system, where the number of subcarriers is set to $N_{s}=1024$ and the subcarrier spacing is $f_{s}=15$ kHz. For the Orth-Com benchmark, the entire bandwidth is uniformly assigned to $K_{\rm T}$ devices; hence, the rate of each device can be approximated as [38] $R_{\text{Orth}}=\dfrac{N_{s}f_{s}}{K_{\rm T}}{\log_{2}}\left(1+\dfrac{1.5}{-\ln\left(5\text{BER}\right)}\text{SNR}\right)$. Then, the communication latency of Orth-Com benchmark can be calculated as $N\log_{2}(Q)/R_{\text{Orth}}$. For the proposed ToDMA framework, the latency can be calculated as $LN/(N_{s}f_{s})$, which has no relationship to $K_{\rm T}$. The numerical comparison at SNR=$25$ dB is shown in Fig. 5. It can be seen that the latency of the proposed ToDMA is several times less than that of the Orth-Com scheme, especially for a larger total number of devices $K_{\rm T}$.

### v2

@@ II-B Evolution of Multiple Access in 6G
Advancements in multiple access towards 6G focus on efficient random access for numerous uncoordinated devices with sporadic activity and short data packets [28, 29, 30]. Key 6G technologies like NOMA and grant-free random access (GFRA) are expected to meet these demands [31]. For example, in GFRA, after receiving a beacon signal from the base station (BS), devices transmit their unique preambles over the same time-frequency resources. The BS then performs device activity detection, channel state information (CSI) estimation, and data detection. This protocol is commonly utilized in studies such as [32, 33, 34]. To accommodate a larger number of devices, unsourced massive access (UMA) introduces a shared preamble codebook, where messages are represented as indices of codewords within the codebook. The BS focuses on decoding the transmitted codewords from the shared codebook, without necessarily associating them to their transmitters [35, 36, 37].
@@ IV-A2 Token Modulation
Assuming that the same tokenizer is shared among the devices, we propose to use a shared modulation codebook for multiple access, denoted as ${\bf U}\in\mathbb{C}^{L\times Q}$, where $L$ denotes the length of each codeword. For simplicity, we consider the columns of ${\bf U}$ are orthonormal vectors, i.e., ${\bf U}^{H}{\bf U}=\mathbf{I}_{Q}$. We use $g(\cdot)$ to represent the token modulation process; hence, the output is denoted by ${\bf X}_{k}={\bf U}{\bf B}_{k}=g({\bf B}_{k})\in\mathbb{C}^{L\times N}$, where each token is modulated to a modulation codeword. We denote the $n$-th column of ${\bf X}_{k}$ as ${\bf x}_{k,n}={\bf U}{\bf b}_{k,n}\in\mathbb{C}^{L}$. The advantages of the proposed token modulation are mainly two fold: ① Shared modulation codebook enables uncoordinated random access of massive number of devices, which reduces the access latency and is easy for implementation. ② The basic communication unit changes from the conventional “bits” to “tokens”, which facilitates applying transformer-based token processing and MLLM s at the receiver.
@@ IV-B Multiple Access Signal Model
$={\bf U}\sum\limits_{k\in[K]}{\bf b}_{k,n}\left({\bf h}_{k}\right)^{T}+{\bf Z}_{n}={\bf U}{\bf H}_{n}+{\bf Z}_{n},\vspace{-2mm}$ (1)
where ${\bf h}_{k}\in\mathbb{C}^{M}$ denotes the channel vector between the $k$-th device and the BS, ${\bf H}_{n}=\sum\nolimits_{k\in[K]}{\bf b}_{k,n}({\bf h}_{k})^{T}\in\mathbb{C}^{Q\times M}$ is termed as an equivalent channel matrix, and the noise matrix ${\bf Z}_{n}$ follows an independent and identically distributed (i.i.d.) complex Gaussian distribution with zero mean and variance $\sigma^{2}{\bf I}_{n}$. Without loss of generality, the channel vector ${\bf h}_{k}$, $\forall k$, is assumed to remain the same throughout the $N$ time slots.
@@ V-C Token Assignment
Here, the CSI acts as a matching anchor that links detected tokens to their respective devices. We assume the CSI of $K$-devices are known, denoted by $\{{\bf h}_{1},{\bf h}_{2},...,{\bf h}_{K}\}$. We initialize the estimated token sequences as ${\bf c}_{k}={\bf 0}_{N}$, $\forall k\in[K]$. Note that the elements of ${\bf c}_{k}$ shall be integers within $[Q]$ to represent tokens, which is equivalent to the one-hot vector representation introduced in Section IV-A. The token assignment process is described as follows. ① Initial Assignment: We define a residual token set $\mathcal{P}_{n}^{r}$, which is initialized as $\mathcal{P}_{n}^{r}=\mathcal{P}_{n}$, $\forall n\in[N]$. We iterate $k$ from $1$ to $K$, in each iteration, we first find the token that minimize the Euclidean distance:
@@
where $\|\cdot\|_{2}$ denotes the $l_{2}$ norm. Next, if and only if $\|{\bf h}_{k}^{0}-{\bf h}_{\phi^{\star}_{n},n}\|_{2}^{2}/M<T_{h}$, we assign token $\phi^{\star}_{n}$ to $[{\bf c}_{k}]_{n}$, and then update $\mathcal{P}_{n}^{r}$ by deleting $\phi^{*}_{n}$; Otherwise, we assign “[MASK]” to $[{\bf c}_{k}]_{n}$. ② Fine-Grained Update: After the initial assignment, if $\mathcal{P}_{n}^{r}=\emptyset$, there is no token collision in this time slot; if $|\mathcal{P}_{n}^{r}|=1$, we assign the only residual token in $\mathcal{P}_{n}^{r}$ to the “[MASK]” positions of $[{\bf c}_{k}]_{n}$, $\forall k\in[K]$; If $|\mathcal{P}_{n}^{r}|>1$, the elements of this residual token set serve as candidate tokens to assist the following steps.
@@ VI-A3 Token Communication Metrics
We define token error rate (TER) as $\text{TER}=\frac{1}{2NK}\sum_{k=1}^{K}\|\widehat{\bf B}_{k}-{\bf B}_{k}\|_{0}.$ It is worth emphasizing that TER is a general metric for ToDMA without specific data modalities. To evaluate the image transmission performance, we consider the well-known PSNR and LPIPS to measure the distortion and perception qualities, respectively.
@@ VI-B Latency Comparisons
To calculate the communication latency, we consider an OFDM system, where the number of subcarriers is set to $N_{s}=1024$ and the subcarrier spacing is $f_{s}=15$ kHz. For the Orth-Com benchmark, the entire bandwidth is uniformly assigned to $K_{\rm T}$ devices; hence, the rate of each device can be approximated as [40] $R_{\text{Orth}}=\frac{N_{s}f_{s}}{K_{\rm T}}{\log_{2}}(1+\frac{1.5}{-\ln\left(5\text{BER}\right)}\text{SNR})$. Then, the communication latency of Orth-Com benchmark can be calculated as $N\log_{2}(Q)/R_{\text{Orth}}$. For the proposed ToDMA framework, the latency can be calculated as $LN/(N_{s}f_{s})$, which has no relationship to $K_{\rm T}$. The numerical comparison at SNR=$25$ dB is shown in Fig. 5. It can be seen that the latency of the proposed ToDMA is several times less than that of the Orth-Com scheme, especially for a larger total number of devices $K_{\rm T}$.